OpenStudy (magbak):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! I WILL AWARD MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! With the restriction x ≠ 2, which of the following is the simplified form of 2x -7 x-5 ----- - ---- x-2 x-2

OpenStudy (magbak):

@agent0smith @zscdragon @Mousam @Mertsj

OpenStudy (magbak):

@agent0smith

OpenStudy (anonymous):

denominator is the same so i guess you just do normal fractional subtraction...

OpenStudy (anonymous):

since due to restrictions the denominator cannot be zero,

OpenStudy (anonymous):

what are the option??

OpenStudy (magbak):

One sec please

OpenStudy (magbak):

1 x-12 ---- x-2 x-2 2x+12 ------ x-2

OpenStudy (anonymous):

it is 1

OpenStudy (mathmate):

@mousam can you explain why?

OpenStudy (anonymous):

\[\frac{ 2x-7 }{ x-2 } - \frac{ x-5 }{ x-2 }\]\[\frac{ x-2 }{ x-2 }\]\[1\]

OpenStudy (magbak):

Ok

OpenStudy (magbak):

Still have more questions please come when tagged.LOL

OpenStudy (anonymous):

lol is this your homework?

OpenStudy (mathmate):

@mousam Thank you!